# Turning Up the Heat: Min-p Sampling for Creative and Coherent LLM Outputs

### v2

@@ Abstract
Large Language Models (LLMs) generate text by sampling the next token from a probability distribution over the vocabulary at each decoding step. However, popular sampling methods like top- (nucleus sampling) often struggle to balance quality and diversity, especially at higher temperatures, leading to incoherent or repetitive outputs. To address this challenge, we propose min- sampling, a dynamic truncation method that adjusts the sampling threshold based on the modelβs confidence by scaling according to the top tokenβs probability. We conduct extensive experiments on benchmarks including GPQA, GSM8K, and AlpacaEval Creative Writing, demonstrating that min- sampling improves both the quality and diversity of generated text, particularly at high temperatures. Moreover, human evaluations reveal a clear preference for min- sampling in terms of both text quality and diversity. Min- sampling has been adopted by multiple open-source LLM implementations, highlighting its practical utility and potential impact.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success in generating coherent and creative text across diverse domains, from factual question answering to open-ended storytelling. A central challenge in these generative tasks is managing the trade-off between creativity and coherence, often influenced by the sampling strategy used during text generation. Popular methods like top- sampling (nucleus sampling) (Holtzman etΒ al., 2020) and temperature scaling (Ackley etΒ al., 1985) are widely adopted to address this challenge, but they often struggle, especially at higher temperatures. While increasing temperature can enhance diversity, it frequently reduces the coherence of generated text; conversely, more conservative sampling limits creativity and can lead to repetitive outputs.
This challenge becomes particularly critical when LLMs are used in tasks that require both imaginative and contextually grounded responses. In this paper, we address this fundamental issue by introducing a new sampling method called min- sampling, designed to dynamically balance creativity and coherence, even at high temperatures. Min- sampling establishes a minimum base probability threshold that scales according to the top tokenβs probability, allowing it to dynamically include diverse options when the model is uncertain while focusing on high-confidence tokens when the model is confident.
@@
We present comprehensive experimental results on several benchmarks, demonstrating that min- consistently improves the quality and diversity of generated text compared to top- sampling and other existing methods.
@@
The rapid adoption of min- by the open-source LLM community further highlights its effectiveness and practical potential.
@@ 2 Related Work
Recent work has introduced methods such as entropy-dependent truncation (-sampling) and mirostat sampling, which attempt to dynamically adjust the sampling pool based on the entropy of the token distribution (Hewitt etΒ al., 2022a; Basu etΒ al., 2021; ShwartzΒ Ziv & LeCun, 2024). While entropy/uncertainty-based approaches show promise in improving text quality, they often require complex parameter tuning and are computationally expensive, making them challenging to use in practical applications. We detail our experimental challenges running sampling in SectionΒ B.2.
@@ 3.1 Overview of Min- Sampling
Formally, at each time step , let denote the vocabulary, and represent the conditional probability distribution over the vocabulary for the next token . Min- sampling involves the following steps:
@@ 3.3 Advantages Over Existing Methods
A primary limitation of existing sampling methods is their performance at high temperatures. As the temperature increases, the token probabilities become more uniform, allowing unlikely tokens to be selected, which can result in incoherent text. Min- addresses this issue by scaling the truncation threshold proportionally to the modelβs confidence, ensuring that the output remains sensible even at higher temperatures. This capability is particularly valuable for tasks that benefit from high creativity, such as storytelling and dialogue generation.
@@
Min- sampling retains computational simplicity, requiring only a few additional calculations over standard top- sampling. Unlike methods that involve auxiliary models or complex entropy-based adjustments, min- can be easily integrated into existing LLM inference pipelines without significant overhead. This makes it practical for both research and real-world applications, and offers a distinct advantage over other entropy-based methods such as and sampling (Hewitt etΒ al., 2022b), as we discuss in SectionΒ B.2.
@@ 3.4 Implementation Details
Min- sampling can be implemented as a logits processor in frameworks like Hugging Face Transformers (Wolf etΒ al., 2020). After applying temperature scaling, the scaled threshold is computed, and tokens with probabilities below this threshold are filtered out before sampling. These operations are efficiently implemented using vectorized computations, adding negligible overhead to the decoding process.
@@
Choosing the Base Threshold (): Setting between 0.05 and 0.1 provides a good balance between creativity and coherence across various tasks and models. Higher values of (e.g., close to 1) can be used to maintain coherence at very high temperatures.
@@
Combining with Other Techniques: While min- sampling can be used in conjunction with other sampling methods or repetition penalties, it is recommended to use it as the primary truncation method to fully leverage its dynamic capabilities.
@@ 3.5 Availability of Resources
Hugging Face Transformers: An implementation is available as a custom logits processor that can be integrated into the generation pipeline.
@@
β’
Project Repository: Code, usage examples, and integration guides are available at our project repository.111https://github.com/menhguin/minp_paper/
β’
Community Adoption: Min- sampling has been rapidly adopted by the open-source community, with over 54,000 GitHub repositories using it, amassing a cumulative 1.1 million stars across these projects.
@@ 5 Experiments
We comprehensively evaluated min- sampling compared to existing methods across multiple benchmarks and model sizes. Our experiments aimed to demonstrate that min- sampling effectively balances creativity and coherence, particularly at higher temperatures.
@@ 5.1 Experimental Setup
Experiments were conducted using the Mistral 7B language model (Jiang etΒ al., 2023), selected for its strong performance across various tasks. To evaluate whether the benefits of min- sampling scale to larger models, we also performed tests on Mistral Large with 123B parameters.
@@
We compared min- sampling against baseline methods, including top- sampling (Holtzman etΒ al., 2020), temperature sampling, sampling (Hewitt etΒ al., 2022b), sampling (Hewitt etΒ al., 2022b) and mirostat sampling (Basu etΒ al., 2021). We present results between temperatures 0.7 and 3.0, with further tests between 0 and 5 linked in our project repository 222https://github.com/menhguin/minp_paper/.
For min-, base probability thresholds of and were used, while top- sampling employed . These hyperparameter settings were chosen based on empirical guidelines and prior research to provide a fair comparison (See SectionΒ B.1 for extensive discussion).
@@ Results
As shown in TableΒ 2, min- consistently outperforms the other methods in almost all temperature settings. The performance advantage becomes more pronounced at higher temperatures, indicating that min- sampling preserves the modelβs problem-solving abilities even when generating more diverse outputs. We also observed significant differences in test-time computing across sampling methods, as detailed in SectionΒ B.2, where and sampling exhibited exponential runtime increases with temperature compared to min-, and failed to load on altogether.
@@
To further understand the accuracy-diversity tradeoff, we evaluate both metrics on the GSM8K dataset using chain-of-thought reasoning with using self-consistency decoding (Wang etΒ al., 2022). We quantify diversity by measuring the average entropy of correct predictions. Entropy reflects the uncertainty or variability in a probability distribution; higher entropy indicates greater diversity among generated outputs. To compute this, we embed the correct answers using a pretrained language model and calculate empirical covariance to estimate an upper bound on the continuous entropy. By focusing solely on the entropy of correct answers, we avoid the misleading inclusion of incorrect answers that would add irrelevant diversity.
The results shown in FigureΒ 2 illustrate that min- sampling achieves a better trade-off between accuracy and creativity compared to top- sampling. Min- sampling consistently lies closer to the Pareto frontier, indicating more efficient performance. The greater spread of min- configurations shows its sensitivity to hyperparameter settings, allowing fine-grained control over the diversity and coherence of outputs, whereas top- configurations cluster strongly, showing that top- sampling is less sensitive to hyperparameter values. We further discuss this nuance in SectionΒ B.3.
@@ 5.2.3 Creative Writing
We used the AlpacaEval Creative Writing benchmark to assess the modelβs ability to generate creative and engaging text (Li etΒ al., 2023). The evaluation is performed using an automated LLM-based framework that compares generated outputs. Similarly to Gusev (2023), we report both the win rate and the length-controlled win rate (LC-Win Rate), which controls for differences in output length.
@@ 5.3 Ablation Study
We conducted an ablation study on the AlpacaEval Creative Writing benchmark to evaluate the impact of different parameter settings on the performance of the min- sampling method. Specifically, we compared min- and top- sampling across various temperatures and configurations. We used two key metrics: Winrate and Winrate (LC).
@@ 6 Human Evaluation
To complement our quantitative benchmarks and explore the qualitative benefits of min- sampling, we conducted a comprehensive human evaluation focusing on creative writing. This evaluation aimed to assess the perceived quality and diversity of text generated using min- and top- sampling at various temperature settings.
@@
To ensure high-quality responses, we implemented several measures:
@@
Attention Checks and Anti-AI Measures: We included attention checks to filter out participants who did not carefully read instructions or samples, and adversarial prompts to filter out automated LLM responses. Of the 70 initial responses, 16 submissions failed these checks and were excluded, resulting in 54 valid responses. We provide an example of our attention check for reference:
@@
| (a) Accuracy (%) on GPQA Main benchmark (Mistral Large) Method Tempβ Only 37.72 31.25 29.02 20.09 2.90 0.89 Top- 39.51 33.26 29.24 18.75 2.01 0.67 Top- 40.18 34.38 29.69 21.21 2.01 0.89 Min- (ours) 38.17 34.60 31.03 27.46 22.77 13.84 | (b) Win rates (%) on AlpacaEval Creative Writing benchmark Method Temperature Only 49.97 53.18 Mirostat 16.69 14.23 Sampling 43.50 45.51 Top- 50.07 β Min- (ours) 52.01 56.54 |
@@
For this creative writing task, we prompted a Llama 3 70B model with the prompt "Write me a creative story?" at various top-, min- and temperature settings. For each sampling method, outputs were generated at three temperature settings: , , and . For each temperature, we tested two diversity levels: low diversity with top- = 0.1 and min- = 0.2, and high diversity with top- = 0.9 and min- = 0.05. Participants were presented with three samples for each configuration to assess output diversity.
@@
Participants rated each set of outputs on two criteria, using a scale from 1 (lowest) to 10 (highest):
@@
Output Quality: Assessed based on how well the outputs fulfilled the prompt, including coherence, relevance, and overall writing quality.
@@
Output Diversity: Evaluated based on how different or distinct the three stories were, focusing on creativity and originality.
@@ Results
TableΒ 4 summarizes the average scores for quality and diversity across different temperature and diversity settings. Overall, min- sampling consistently scored higher than top- sampling across all settings. At higher temperatures, while top- samplingβs scores for quality and diversity decreased significantly, min- sampling maintained high scores. A paired t-test confirmed that the differences in scores between min- and top- sampling were statistically significant (p < 0.05).
Qualitative Feedback
Participants frequently noted that outputs generated with min- sampling were more coherent and creative, especially at higher temperatures. In contrast, top- sampling often produced incoherent, less diverse outputs in similar conditions, especially with low diversity settings.
| Temperature | Diversity Setting | Min- | Top- | ||
| Quality | Diversity | Quality | Diversity | ||
| 1.0 | Low | 7.06 1.48 | 5.83 2.03 | 5.96 2.24 | 2.40 2.01 |
| High | 8.02 1.35 | 7.74 1.63 | 7.67 1.38 | 7.04 1.88 | |
| 2.0 | Low | 7.62 1.53 | 6.91 1.94 | 5.43 2.24 | 1.83 1.61 |
| High | 7.98 1.42 | 7.96 1.54 | 7.75 1.37 | 7.66 1.50 | |
| 3.0 | Low | 7.74 1.76 | 7.60 1.86 | 5.75 2.33 | 2.25 2.44 |
| High | 7.57 1.68 | 7.66 1.45 | 7.11 2.09 | 7.49 1.74 | |

### v4

# Turning Up the Heat: Min- Sampling for Creative and Coherent LLM Outputs

@@ Abstract
Large Language Models (LLMs) generate text by sampling the next token from a probability distribution over the vocabulary at each decoding step. Popular sampling methods like top- (nucleus sampling) often struggle to balance quality and diversity, especially at higher temperatures which lead to incoherent or repetitive outputs. We propose min- sampling, a dynamic truncation method that adjusts the sampling threshold based on the modelβs confidence by using the top tokenβs probability as a scaling factor. Our experiments on benchmarks including GPQA, GSM8K, and AlpacaEval Creative Writing show that min- sampling improves both the quality and diversity of generated text across different model families (Mistral and Llama 3) and model sizes (1B to 123B parameters), especially at higher temperatures. Human evaluations further show a clear preference for min- sampling, in both text quality and creativity. Min- sampling has been adopted by popular open-source LLM frameworks, including Hugging Face Transformers, VLLM, and many others, highlighting its significant impact on improving text generation quality.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success in generating coherent and creative text across diverse domains, from factual question answering to open-ended storytelling. A central challenge in these generative tasks is managing the trade-off between creativity and coherence, often influenced by the sampling strategy used during text generation. Popular methods like Top- sampling (nucleus sampling) (Holtzman etΒ al., 2020) and temperature scaling (Ackley etΒ al., 1985) are widely adopted to address this challenge. While increasing temperature can enhance diversity, it frequently reduces the coherence of generated text, even with additional Top- sampling. Conversely, more conservative sampling limits creativity and can lead to repetitive outputs.
In this paper, we address this fundamental issue by introducing a new sampling method called min- sampling, designed to dynamically balance creativity and coherence, even at high temperatures. Min- sampling establishes a minimum base probability threshold that scales according to the top tokenβs probability, allowing it to dynamically include diverse options when the model is uncertain while focusing on high-confidence tokens when the model is confident.
@@
We present comprehensive experimental results across several benchmarks, model families and model sizes, demonstrating that min- consistently improves the quality and diversity of generated text compared to Top- sampling and other existing methods.
@@
The rapid adoption of min- by the open-source LLM community highlights its practical benefits, even for large reasoning models such as Deepseek R1 (DeepSeek-AI etΒ al., 2025).
@@ 2 Related Work
Recent work has introduced methods such as entropy-dependent truncation (-sampling) and mirostat sampling, which attempt to dynamically adjust the sampling pool based on the entropy of the token distribution (Hewitt etΒ al., 2022a; Basu etΒ al., 2021). While entropy/uncertainty-based approaches show promise in improving text quality, they often require complex parameter tuning and are computationally expensive, making them challenging to use in practical applications. We detail our experimental challenges running sampling in SectionΒ A.3.
@@ 3.1 Overview of Min- Sampling
Formally, at each time step , let denote the vocabulary, and represent the conditional probability distribution over the vocabulary for the next token . Min- sampling works as follows:
@@ 3.3 Advantages Over Existing Methods
A primary limitation of existing sampling methods is their performance at high temperatures. As the temperature increases, the token probabilities become more uniform, allowing unlikely tokens to be selected, which can result in incoherent text. Min- addresses this issue by scaling the truncation threshold proportionally to the modelβs confidence, ensuring that the output remains sensible even at higher temperatures, making it increasingly popular for tasks that benefit from high creativity, such as storytelling or generating diverse solutions (Tajwar etΒ al., 2025).
@@
Min- sampling retains computational simplicity, requiring only a few additional calculations over standard Top- sampling. Unlike methods that involve auxiliary models or complex entropy-based adjustments, min- can be easily integrated into existing LLM inference pipelines without significant overhead. This makes it practical for both research and real-world applications, and offers a distinct advantage over other entropy-based methods such as and sampling (Hewitt etΒ al., 2022b), as we discuss in SectionΒ A.3.
@@ 3.4 Implementation Details
Min- sampling can be implemented as a logits processor in frameworks like Hugging Face Transformers (Wolf etΒ al., 2020) and , vLLM (Kwon etΒ al., 2023). After applying temperature scaling, the scaled threshold is computed, and tokens with probabilities below this threshold are filtered out before sampling. These operations are efficiently implemented using vectorized computations, adding negligible overhead to the decoding process.
@@
Choosing the Base Threshold (): Setting between 0.05 and 0.1 provides a good balance between creativity and coherence across various tasks and models. Higher values of (e.g., close to 1) can be used to maintain coherence at very high temperatures. Higher values of (e.g., 0.5-0.7) can further increase coherence and benchmark scores at high temperatures, though likely at the cost of diversity (see TableΒ 13 for detailed analysis).
@@
Combining with Other Techniques: While min- sampling can be used in conjunction with other sampling methods or repetition penalties, it is recommended to use it as the primary truncation method to fully leverage its dynamic capabilities. Our experiments with combined samplers (see TableΒ 12) show that using multiple truncation methods can lead to double normalization issues and suboptimal performance, as each methodβs hyperparameters have so far only been optimised for standalone application. We discuss potential future research directions for combined sampling approaches in SectionΒ D.2.
@@ 3.5 Availability of Resources
Community Adoption: Due to popular demand from open-source developers, Min- sampling has seen rapid uptake among high-traffic, production-grade inference engines. It is now integrated in widely used frameworks such as Hugging Face Transformers, vLLM, and SGLang (Zheng etΒ al., 2024), which collectively have accrued over 350,000 GitHub stars. This integration, coupled with extensive downstream usage (e.g., over 290,000 dependent repositories for Transformers alone), underscores the methodβs practical impact (see AppendixΒ A.5 for detailed statistics).
@@
Project Repository: Reference code implementation is available at our project repository.111https://github.com/menhguin/minp_paper
@@
Relevance to advanced reasoning models.
Further validating min-βs practical utility, Unsloth, a major inference provider whose DeepSeek R1 (DeepSeek-AI etΒ al., 2025) implementation has received over 3.5 million downloads as of March 2025, explicitly recommends min- in their official documentation: "Itβs best to use βMin- 0.05 to counteract very rare token predictions" (Unsloth, 2025). This recommendation demonstrates min-βs strong real-world applications.
@@ 5 Experiments
We compared min- with existing sampling methods across benchmarks and models to demonstrate its effectiveness in balancing creativity and coherence, particularly at higher temperatures.
@@ 5.1 Experimental Setup
Our main experiments were conducted using the Mistral 7B language model (Jiang etΒ al., 2023), selected for its strong performance across various tasks. We additionally conduct evaluations on Llama series models (Grattafiori etΒ al., 2024). To evaluate if the benefits of min- sampling scale to larger models, we also perform tests on Mistral Large with 123B parameters and Llama 3.2 70B.
@@
We compared min- sampling against baseline methods, including Top- sampling (Holtzman etΒ al., 2020), temperature sampling, sampling (Hewitt etΒ al., 2022b), sampling (Hewitt etΒ al., 2022b) and mirostat sampling (Basu etΒ al., 2021). We present results between temperatures 0.7 and 3.0, with further tests between 0 and 5 linked in our project repository 222https://github.com/menhguin/minp_paper/. Min- and Top- perform comparably lower temperatures 0 to 0.5 (see TableΒ 10 in Appendix), with differences within error margins.
For min-, base probability thresholds of and were used, while Top- sampling employed . These hyperparameter settings were chosen based on empirical guidelines and prior research to provide a fair comparison (See SectionΒ A.2 for extensive discussion).
@@ Results
Results Across Model Families
Our extensive experiments on Llama models (see TableΒ 8 in Appendix) demonstrate that the benefits of min- sampling are not limited to Mistral models but generalize well across different model families and scales, from 1B to 70B parameters.
@@
TableΒ 2 shows min- outperforming other methods across most temperature settings. At lower temperatures , see TableΒ 10 in Appendix), min- and top- perform similarly within statistical error margins, as decreasing temperature reduces the impact of all truncation methods by shrinking the token candidate pool. Min-βs advantage becomes more significant at higher temperatures, preserving problem-solving abilities even with increased diversity. Additionally, and sampling showed exponential runtime increases with temperature and failed completely at , while min- remained computationally efficient (SectionΒ A.3).
Notably, across all experiments, best scores were achieved with min- across a range of temperatures, rather than greedy decoding, challenging assumptions that greedy decoding is strictly optimal for task benchmarks (Castel etΒ al., 2022; Wiher etΒ al., 2022). See comparisons in TableΒ 10.
@@
To further understand the accuracy-diversity tradeoff, we evaluate both metrics on the GSM8K dataset using chain-of-thought reasoning with using self-consistency decoding (Wang etΒ al., 2022). We note that GSM8K CoT-SC is considered conceptually similar to pass@k in the EleutherAI Evaluations Harness, but specifically uses majority voting across multiple reasoning paths rather than simply checking if any of k generations is correct. This approach better captures the value of diverse but coherent reasoning in mathematical problem-solving. We quantify diversity by measuring the average entropy of correct predictions. Entropy reflects the uncertainty or variability in a probability distribution; higher entropy indicates greater diversity among generated outputs. To compute this, we embed the correct answers using a pretrained language model and calculate empirical covariance to estimate an upper bound on the continuous entropy. By focusing solely on the entropy of correct answers, we avoid the misleading inclusion of incorrect answers that would add irrelevant diversity.
The results shown in FigureΒ 2 illustrate that min- sampling achieves a better trade-off between accuracy and creativity compared to Top- sampling. Min- sampling consistently lies closer to the Pareto frontier, indicating more efficient performance. The greater spread of min- configurations shows its sensitivity to hyperparameter settings, allowing fine-grained control over the diversity and coherence of outputs, whereas Top- configurations cluster strongly, showing that Top- sampling is less sensitive to hyperparameter values. We further discuss this nuance in SectionΒ A.4.
@@ 5.2.3 Creative Writing
We used the AlpacaEval Creative Writing benchmark to assess the modelβs ability to generate creative and engaging text (Li etΒ al., 2023). This evaluation employs LLM-as-a-judge on generated outputs. Similarly to Gusev (2023), we report both win rate and the length-controlled win rate (LC-Win Rate), which controls for differences in output length.
Results
@@
Independent evaluations on EQ-Bench further support these findings, where min- with temperature 1.5 and scored 62 versus the baselineβs 51.5 at temperature 1.0 (Gusev, 2024). A supplementary structured LLM-as-judge evaluation focusing on specific creative writing dimensions further confirmed min-βs advantages across temperature settings (see SectionΒ C.3.6).
@@ 5.3 Ablation Study
We conducted an ablation study on the AlpacaEval Creative Writing benchmark comparing min- and Top- sampling across different temperatures and parameter configurations. Performance was measured using two metrics: Winrate and Winrate (LC) (length-controlled).
@@ 6 Human Evaluation
To complement our quantitative benchmarks of min- sampling, we conducted a comprehensive human evaluation focusing on creative writing. This evaluation assesses the perceived quality and diversity of text generated using min- and Top- sampling at various temperature settings.
@@
To ensure high-quality responses, we implemented several measures. For full details, see sectionΒ C.1
@@
Attention Checks and Anti-AI Measures: We included attention checks to filter out participants who did not carefully read instructions or samples, and adversarial prompts to filter LLM responses. Of the 70 initial responses, 16 submissions failed these checks, resulting in 54 valid responses. We provide an example of our attention check below:
@@
| (a) Accuracy (%) on GPQA Main benchmark (Mistral Large) Method Tempβ Only 37.72 31.25 29.02 20.09 2.90 0.89 Top- 39.51 33.26 29.24 18.75 2.01 0.67 Top- 40.18 34.38 29.69 21.21 2.01 0.89 Min- (ours) 38.17 34.60 31.03 27.46 22.77 13.84 | (b) AlpacaEval Creative Writing (% win) Method Temperature Only 49.97 53.18 Mirostat 16.69 14.23 Sampling 43.50 45.51 Top- 50.07 β Min- (ours) 52.01 56.54 |
@@
We evaluated creative writing performance using a Llama 3 70B model across different sampling configurations. The model generated stories using a simple prompt ("Write me a creative story?") with either Top- or min-. We tested three temperature settings: and two diversity levels: low (Top- = 0.1 and min- = 0.2), and high (Top- = 0.9, min- = 0.05). This yielded 12 total configurations: 2 sampling methods 3 temperatures 2 diversity settings. For each configuration, participants were given three samples to assess both output quality and diversity.
@@
Participants rated sets of outputs on two criteria from 1 (lowest) to 10 (highest):
@@
Quality: How well the outputs fulfilled the prompt; coherence, relevance, overall quality.
@@
Diversity: How creative or distinct the three stories were.
@@ Results
TableΒ 4 summarizes average quality and diversity scores across temperature and diversity settings. Overall, min- sampling consistently outperformed Top- sampling, especially at higher temperatures where Top-βs quality and diversity scores declined sharply, while min- maintained high scores. To address limitations of the initial evaluation, we conducted a refined human study with an improved inference engine and rigorous setup (detailed in SectionΒ C.2), which reinforced our findings. Both evaluation results are available on our GitHub repository333https://github.com/menhguin/minp_paper/.
| Low Diversity Settings | High Diversity Settings | ||||||
| Temp | Metric | Standard | Top- | Min- | Standard | Top- | Min- |
| Sampling | = 0.1 | = 0.2 | Sampling | = 0.9 | = 0.05 | ||
| T=1 | Quality | 7.550.2365 | 5.960.3077 | 7.060.2039 | 7.960.1532 | 7.670.1931 | 8.020.1858 |
| Diversity | 7.080.2921 | 2.400.2765 | 5.830.2795 | 7.660.2014 | 7.040.2587 | 7.740.2235 | |
| T=2 | Quality | 7.760.2640 | 5.430.3083 | 7.620.2105 | 7.850.1931 | 7.750.1885 | 7.980.1952 |
| Diversity | 7.660.2459 | 1.830.2212 | 6.910.2658 | 7.570.2058 | 7.660.2066 | 7.960.2117 | |
| T=3 | Quality | 6.750.3509 | 5.750.3201 | 7.740.2418 | 6.830.2895 | 7.110.2869 | 7.570.2303 |
| Diversity | 7.040.2836 | 2.250.3353 | 7.600.2550 | 7.430.2830 | 7.490.2396 | 7.660.1996 | |